OpenStudy (anonymous):

Prove the identity: tan x = (cos x - cos 3x + sin 3x - sin x)/(cos x + cos 3x + sin 3x + sin x)

OpenStudy (anonymous):

do u know cos x +cosy =2 cos(x+y)/2 cos(x-y)/2 and other formulas of this form

OpenStudy (anonymous):

see cos x - cos 3x = 2 (sin (x+3x)/2 sin(3x-x)/2) = 2sin2xsinx cos 3x + cos x = 2 (cos (x+3x)/2 cos(3x-x)/2) = 2cos2xcosx sin 3x - sin x = 2(cos (x+3x)/2 sin(3x-x)/2 ) = 2 cos2x sinx sin 3x + sin x = 2(sin (x+3x)/2cos(3x-x)/2 ) = 2 sin2x cosx i hope u will be able to proceed nwo..

OpenStudy (anonymous):

Yes. Thank you!

OpenStudy (anonymous):

yw